ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 107, 1998

Argyros, S. A., S. Mercourakis and A. Tsarpalias Convex unconditionality and summability of weakly null sequences

157

Ball, K. and I. Perissinaki The subindependence of coordinate slabs in l_{p}^{n} balls

289

Casacuberta, C., J. L. Rodríguez and L. J. Hernández Models for torsion homotopy types

301

Cedó, F. and L. H. Rowen Addendum to “Examples of semiperfect rings”

343

Dixon, M. R., M. J. Evans and H. Smith On groups that are residually of finite rank

1

Domínguez, D. A finiteness theorem for transitive foliations and flat vector bundles

253

Evans, M. J. See M. R. Dixon

Hájek, P. Smooth functions on C(K)  237

Heicklen, D. Bernoullis are standard when entropy is not an obstruction

141

Herfort, W. N., L. Ribes and P. A. Zalesskii Finite extensions of free pro-p groups of rank at most two

195

Hernández, L. J. See C. Casacuberta

Herwig, B. Extending partial isomorphisms for the small index property of many ω-categorical structures

93

Lusky, W. Three space properties and basis extensions  17

Margulis, G. A. and S. Mozes Aperiodic tilings of the hyperbolic plane by convex polygons

319

Mercourakis, S. See S. A. Argyros

Mestre, J-F. Extensions Q-régulières de Q(t) de groupe de Galois

6.A_{6} et 6.A_{7} 333

Mozes, S. See G. A. Margulis

Perissinaki, I. See K. Ball

Ribes, L. See W. N. Herfort

Rodríguez, J. L. See C. Casacuberta

Rowen, L. H. See F. Cedó Smith, H. See M. R. Dixon

Smorodinsky, M. Processes with no standard extension  327

Stollmann, P. Localization for random perturbations of anisotropic periodic media

125

Szigeti, J. On the characteristic polynomial of supermatrices  229
Tadić, M. On reducibility of parabolic induction  29

Tsarpalias, A. See S. A. Argyros

Zalesskii, P. A. See W. N. Herfort